# On the Role of Attention Heads in Large Language Model Safety

### v1

@@ 1 Introduction
The capabilities of large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Dubey et al., 2024; Yang et al., 2024) have recently improved significantly while learning from larger pre-training datasets. Despite this, language models may respond to harmful queries, generating unsafe and toxic content (Ousidhoum et al., 2021; Deshpande et al., 2023), raising concerns about potential risks (Bengio et al., 2024). In sight of this, alignment (Ouyang et al., 2022; Bai et al., 2022a; b) is employed to ensure LLM safety by aligning with human values, while existing research (Zou et al., 2023b; Wei et al., 2024a; Carlini et al., 2024) suggests that malicious attackers can circumvent safety guardrails. Therefore, understanding the inner workings of language models is necessary for responsible and ethical development (Zhao et al., 2024a; Bereska & Gavves, 2024).
Currently, revealing the black box of LLM safety mechanisms is typically achieved through mechanism interpretation methods. Specifically, these methods (Geiger et al., 2021; Stolfo et al., 2023; Gurnee et al., 2023) granularly analyze features, neurons, layers and parameters to assist human understand model behavior and capabilities. Recent studies (Zou et al., 2023a; Templeton, 2024; Arditi et al., 2024; Chen et al., 2024) indicates that the safety capability of LLM can be attributed to representations and neurons. However, multi-head attention, which has been confirmed to be crucial in other abilities (Vig, 2019; Gould et al., 2024; Wu et al., 2024), has received less attention in safety interpretability. Due to the differing specificities of components and representations, directly transferring existing methods to the safety attention mechanism attribution task is challenging. Additionally, some general approaches (Meng et al., 2022; Wang et al., 2023; Zhang & Nanda, 2024) typically involve special tasks in order to observe the result changes in one forward, whereas safety tasks necessitate full generation across multiple forwards.
@@
In this paper, we aim to interpret safety capability within multi-head attention. To achieve this, we introduce Safety Head ImPortant Scores (Ships) to attribute the safety capability of individual attention head in an aligned model. Since the model is trained to reject harmful queries with high probability to align with human values (Ganguli et al., 2022; Dubey et al., 2024). Base on this, Ships quantifies the impact of each attention head on the change in the rejection probability of harmful queries through causal tracing. Concretely, we demonstrate that Ships can be used for attributing safety attention head. Experimental results show that on three harmful query datasets, using Ships to identify safe heads and using undifferentiated attention ablation (only modifying $\sim$ 0.006% of the parameters) can improve the attack success rate (ASR) of Llama-2-7b-chat from 0.04 to 0.64 $\uparrow$ and Vicuna-7b-v1.5 from 0.27 to 0.55 $\uparrow$.
Furthermore, to attribute generalized safety attention heads, we generalize Ships to evaluate the changes in the representation of ablating attention heads on harmful query datasets. Based on the generalized version of Ships, we attribute the most important safety attention head, which is ablated, and the ASR is improved to 0.72 $\uparrow$. Iteratively selecting important heads result in a group of heads that can significantly change the rejection representation. We name this heuristic method Safety Attention Head AttRibution Algorithm (Sahara). Experimental results show that ablating the attention head group can further weaken the safety capability collaboratively.
Based on the Ships and Sahara, we interpret the safety head of attention on several popular LLMs, such as Llama-2-7b-chat and Vicuna-7b-v1.5. This interpretation yields several intriguing insights: 1. Certain safety heads within the attention mechanism are crucial for feature integration in safety tasks. Specifically, modifying the value of the attention weight matrices changes the model output significantly, while scaling the attention output does not; 2. For LLMs fine-tuned from the same base model, their safety heads have overlap, indicating that in addition to alignment, the safety impact of the base model is critical; 3. The attention heads that affect safety can act independently with affecting helpfulness little. These insights provide a new perspective on LLM safety and provide a solid basis for the enhancement and future optimization of safety alignment.
Our contributions are summarized as follows:
@@
We present a novel metric, Ships, to evaluate the impact of attention head ablation on safety. Base on this, we propose a heuristic algorithm, termed Sahara, to find attention head groups whose ablation leads to safety degradation.
@@ 2 Preliminary
Large Language Models (LLMs). Current state-of-the-art LLMs are predominantly based on a decoder-only architecture, which predict the next token for the given prompt. For the input sequence $x=x_{1},x_{2},\ldots,x_{s}$, LLMs can return the probability distribution of the next token:
@@
where $o_{s}$ is the last residual stream, and $W$ is the linear function, which mapping $o_{s}$ to the the logits associated with each token in the vocabulary $V$. Sampling from the probability distribution yields a new token $x_{n+1}$. Iterating this process allows to obtain a response $R=x_{s+1},x_{s+2},\ldots,x_{s+R}$.
Multi-Head Attention (MHA). The attention mechanism (Vaswani, 2017) in LLMs plays is critical for capturing the features of input sequence. Prior works (Htut et al., 2019; Clark et al., 2019; Campbell et al., 2023; Wu et al., 2024) demonstrate that individual heads in MHA contribute distinctively across various language tasks. MHA, with $n$ heads, is formulated as follows:
@@
LLM Safety and Jailbreak Attack. LLMs may generate content that is unethical or illegal, raising significant safety concerns. To address the risks, safety alignment (Bai et al., 2022a; Dai et al., 2024) is typically implemented to prevent models from responding to harmful queries $x_{\mathcal{H}}$. Specifically, safety alignment train LLMs $\theta$ to optimize the following objective:
@@
where $D$ is a safety discriminator that flags $R$ as harmful when $D(R)=\operatorname{True}$. Prior studies (Zou et al., 2023b; Liao & Sun, 2024; Jia et al., 2024) show that shifting the probability distribution towards affirmative tokens can significantly improve the attack success rate. Suppressing rejection tokens (Shen et al., 2023; Wei et al., 2024a) yields similar results. These insights highlight that LLM safety relies on maximizing the probability of generating rejection tokens in response to harmful queries.
Safety Parameters. Mechanistic interpretability (Zhao et al., 2024a; Lindner et al., 2024) attributes model capabilities to specific parameters, improving the transparency of black-box LLMs while addressing concerns about their behavior. Recent work (Wei et al., 2024b; Chen et al., 2024) specialize in safety by identifying critical parameters responsible for ensuring LLM safety. When these safety-related parameters are modified, the safety guardrails of LLMs are compromised, potentially leading to the generation of unethical content. Consequently, safety parameters are those whose ablation results in the significantly increase in the probability of generating an illegal or unethical response to the harmful queries $x_{\mathcal{H}}$. Formally, we define the Safety Parameters as:
@@
$\Delta p(\theta_{\mathcal{C}})$ $=p\left(R_{\bot}\mid x_{\mathcal{H}};\theta_{\mathcal{O}}\right)-p\left(R_{\bot}\mid x_{\mathcal{H}};(\theta_{\mathcal{O}}\setminus\theta_{\mathcal{C}})\right),$
@@ 3 Safety Head ImPortant Score
In this section, we aim to identify the safety parameters within the multi-head attention mechanisms for a specific harmful query. In Section 3.1, we detail two modifications to ablate the specific attention head for the harmful query. Base on this, Section 3.2 introduces Ships, a method to attribute safety parameter at the head-level based on attention head ablation. Finally, the experimental results in Section 3.3 demonstrate the effectiveness of our attribution method.
@@ 3.1 Attention Head Ablation
We focus on identifying the safety parameters within attention head. Prior studies (Michel et al., 2019; Olsson et al., 2022; Wang et al., 2023) have typically employed ablation techniques by directly setting the attention head outputs to 00. The resulting modified multi-head attention can be formalized as:
$\operatorname{MHA}^{\mathcal{A}}_{W_{q},W_{k},W_{v}}=(h_{1}\oplus h_{2}\cdots\oplus h^{m}_{i}\cdots\oplus h_{n})W_{o},$ (6)
@@
Undifferentiated Attention. Specifically, scaling $W_{q}$ or $W_{k}$ matrix forces the attention weights of the head to collapse to the mean $d_{k}/n$. Note that modifying either $W_{q}$ or $W_{k}$ has equivalent effects, a derivation is given in Appendix A.1. Undifferentiated Attention achieves ablation by hindering the head to extract the critical information from the input sequence. Formally, it can be expressed as:
$h_{i}^{m}$ $=\operatorname{Softmax}\Big{(}\frac{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\epsilon}W_{q}^{i}W_{k}^{i}{}^{T}}{\sqrt{d_{k}/n}}\Big{)}W_{v}^{i}=\frac{n}{d_{k}}W_{v}^{i}.$ (7)
Scaling Contribution. This method scales the attention head output by multiplying $W_{v}$ by $\epsilon$. When the outputs of all heads are concatenated and then multiplied by the fully connected matrix $W_{o}$, the contribution of the modified head $h_{i}^{m}$ is significantly diminished compared to the others. A detailed discussion of scaling the $W_{v}$ matrix can be found in Appendix A.2. This method is similar in form to Undifferentiated Attention, and expressed as:
$h_{i}^{m}$ $=\operatorname{Softmax}\Big{(}\frac{W_{q}^{i}W_{k}^{i}{}^{T}}{\sqrt{d_{k}/n}}\Big{)}{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\epsilon}W_{v}^{i}.$ (8)
@@ 3.2 Evaluate the Importance of Parameters for Specific Harmful Query
Previous studies (Wang et al., 2024; Zhou et al., 2024) find rejection responses to various harmful queries are highly consistent. Furthermore, modern language models tend to be sparse, with many redundant parameters (Frantar & Alistarh, 2023; Sun et al., 2024a; b), meaning ablating some heads often has minimal impact on overall performance. Therefore, when a head is ablated, any deviation from the original rejection distribution suggests a shift towards affirmative responses, indicating that the ablated head is most likely safety parameter.
@@ 3.3 Ablate Attention Heads For Specific Query Impact Safety
Figure 2: Attack success rate (ASR) for harmful queries after ablating the most important safety attention head (bars with x-axis labels ‘Greedy’ and ‘Top-5’), calculated using Ships. And ‘template’ means using chat template as input, ‘direct’ means direct input (refer to Appendix B.2 for detailed introduce). Figure 2(a) shows results with undifferentiated attention, while Figure 2(b) uses scaling contribution.
We conduct a preliminary experiments to demonstrate that Ships can be used to effectively identify safety heads. Our experiments are performed on two models, i.e., Llama-2-7b-chat (Touvron et al., 2023) and Vicuna-7b-v1.5 (Zheng et al., 2024b), using three commonly used harmful query datasets: Advbench (Zou et al., 2023b), Jailbreakbench (Chao et al., 2024), and Malicious Instruct (Huang et al., 2024). After ablating the safety attention head for the specific $q_{\mathcal{H}}$, we generate an output of 128 tokens for each query to evaluate the impact on model safety. We use greedy sampling to ensure result reproducibility and top-k sampling to capture changes in the probability distributions. We use the attack success rate (ASR) metric which is widely used to evaluate model safety (Qi et al., 2024; Zeng et al., 2024):
@@
where $Q_{\text{harm}}$ denotes a harmful query dataset. A higher ASR implicates that the model is more susceptible to attacks thus less safe. The results in Figure 2, indicates that ablating the attention head with the highest Ships score significantly reduces the safety capability. For Llama-2-7b-chat, using undifferentiated attention with chat template, ablating the most important head (which constitutes 0.006% of all parameters) improves the average ASR from 0.04 to 0.64 $\uparrow$ for ‘template’, representing a 16x $\uparrow$ improvement. For Vicuna-7b-v1.5, the improvement is less pronounced but still notable, with an observed improvement from 0.27 to 0.55 $\uparrow$. In both models, Undifferentiated Attention consistently outperforms Scaling Contribution in terms of its impact on safety.
Takeaway. Our experimental results demonstrate that the special attention head can significantly impact on safety in language models, as captured by our proposed Ships metric.
@@ 4 Safety Attention Head AttRibution Algorithm
Building on these insights, we collect the activations $a$ of the top layer across the dataset. We stack the $a$ of all harmful queries into a matrix $M$ and apply SVD decomposition to it, aiming to analyze the impact of ablating attention heads at the dataset level. The SVD of $M$ is expressed as $\operatorname{SVD}(M)=U\Sigma V^{T}$, where the left singular matrix $U_{\theta}$ is an orthogonal matrix of dimensions $\mid Q_{\mathcal{H}}\mid\times d$, representing key feature in the representations space of the harmful query dataset $Q_{\mathcal{H}}$.

### v2

@@ 1 Introduction
The capabilities of large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Dubey et al., 2024; Yang et al., 2024) have significantly improved while learning from larger pre-training datasets recently. Despite this, language models may respond to harmful queries, generating unsafe and toxic content (Ousidhoum et al., 2021; Deshpande et al., 2023), raising concerns about potential risks (Bengio et al., 2024). In sight of this, alignment (Ouyang et al., 2022; Bai et al., 2022a; b) is employed to ensure LLM safety by aligning with human values, while existing research (Zou et al., 2023b; Wei et al., 2024a; Carlini et al., 2024) suggests that malicious attackers can circumvent safety guardrails. Therefore, understanding the inner workings of LLMs is necessary for responsible and ethical development (Zhao et al., 2024a; Bereska & Gavves, 2024; Fang et al., 2024).
Currently, revealing the black-box LLM safety is typically achieved through mechanism interpretation methods. Specifically, these methods (Geiger et al., 2021; Stolfo et al., 2023; Gurnee et al., 2023) granularly analyze features, neurons, layers, and parameters to assist humans in understanding model behavior and capabilities. Recent studies (Zou et al., 2023a; Templeton, 2024; Arditi et al., 2024; Chen et al., 2024) indicate that the safety capability can be attributed to representations and neurons. However, multi-head attention, which is confirmed to be crucial in other abilities (Vig, 2019; Gould et al., 2024; Wu et al., 2024), has received less attention in safety interpretability. Due to the differing specificities of components and representations, directly transferring existing methods to safety attention attribution is challenging. Additionally, some general approaches (Meng et al., 2022; Wang et al., 2023; Zhang & Nanda, 2024) typically involve special tasks to observe the result changes in one forward, whereas safety tasks necessitate full generation across multiple forwards.
@@
In this paper, we aim to interpret safety capability within multi-head attention. To achieve this, we introduce Safety Head ImPortant Scores (Ships) to attribute the safety capability of individual attention heads in an aligned model. The model is trained to reject harmful queries in a high probability so that it aligns with human values (Ganguli et al., 2022; Dubey et al., 2024). Based on this, Ships quantifies the impact of each attention head on the change in the rejection probability of harmful queries through causal tracing. Concretely, we demonstrate that Ships can be used for attributing safety attention head. Experimental results show that on three harmful query datasets, using Ships to identify safe heads and using undifferentiated attention ablation (only modifying $\sim$ 0.006% of the parameters) can improve the attack success rate (ASR) of Llama-2-7b-chat from 0.04 to 0.64 $\uparrow$ and Vicuna-7b-v1.5 from 0.27 to 0.55 $\uparrow$.
Furthermore, to attribute generalized safety attention heads, we generalize Ships to evaluate the changes in the representation of ablating attention heads on harmful query datasets. Based on the generalized version of Ships, we attribute the most important safety attention head, which is ablated, and the ASR is improved to 0.72 $\uparrow$. Iteratively selecting important heads results in a group of heads that can significantly change the rejection representation. We name this heuristic method Safety Attention Head AttRibution Algorithm (Sahara). Experimental results show that ablating the attention head group can further weaken the safety capability collaboratively.
Based on the Ships and Sahara, we interpret the safety head of attention on several popular LLMs, such as Llama-2-7b-chat and Vicuna-7b-v1.5. This interpretation yields several intriguing insights: 1. Certain safety heads within the attention mechanism are crucial for feature integration in safety tasks. Specifically, modifying the value of the attention weight matrices changes the model output significantly, while scaling the attention output does not; 2. For LLMs fine-tuned from the same base model, their safety heads have overlap, indicating that in addition to alignment, the safety impact of the base model is critical; 3. The attention heads that affect safety can act independently with affecting helpfulness little. These insights provide a new perspective on LLM safety and provide a solid basis for the enhancement and future optimization of safety alignment. Our contributions are summarized as follows:
@@
We present Ships to evaluate the safety impact of attention head ablation. Then, we propose a heuristic algorithm Sahara to find head groups whose ablation leads to safety degradation.
@@ 2 Preliminary
Large Language Models (LLMs). Current state-of-the-art LLMs are predominantly based on a decoder-only architecture, which predicts the next token for the given prompt. For the input sequence $x=x_{1},x_{2},\ldots,x_{s}$, LLMs can return the probability distribution of the next token:
@@
where $o_{s}$ is the last residual stream, and $W$ is the linear function, which maps $o_{s}$ to the the logits associated with each token in the vocabulary $V$. Sampling from the probability distribution yields a new token $x_{n+1}$. Iterating this process allows to obtain a response $R=x_{s+1},x_{s+2},\ldots,x_{s+R}$.
Multi-Head Attention (MHA). The attention mechanism (Vaswani, 2017) in LLMs plays is critical for capturing the features of the input sequence. Prior works (Htut et al., 2019; Clark et al., 2019b; Campbell et al., 2023; Wu et al., 2024) demonstrate that individual heads in MHA contribute distinctively across various language tasks. MHA, with $n$ heads, is formulated as follows:
@@
LLM Safety and Jailbreak Attack. LLMs may generate content that is unethical or illegal, raising significant safety concerns. To address the risks, safety alignment (Bai et al., 2022a; Dai et al., 2024) is implemented to prevent models from responding to harmful queries $x_{\mathcal{H}}$. Specifically, safety alignment train LLMs $\theta$ to optimize the following objective:
@@
where $D$ is a safety discriminator that flags $R$ as harmful when $D(R)=\operatorname{True}$. Prior studies (Liao & Sun, 2024; Jia et al., 2024) show that shifting the probability distribution towards affirmative tokens can significantly improve the attack success rate. Suppressing rejection tokens (Shen et al., 2023; Wei et al., 2024a) yields similar results. These insights highlight that LLM safety relies on maximizing the probability of generating rejection tokens in response to harmful queries.
Safety Parameters. Mechanistic interpretability (Zhao et al., 2024a; Lindner et al., 2024) attributes model capabilities to specific parameters, improving the transparency of black-box LLMs while addressing concerns about their behavior. Recent work (Wei et al., 2024b; Chen et al., 2024) specializes in safety by identifying critical parameters responsible for ensuring LLM safety. When these safety-related parameters are modified, the safety guardrails of LLMs are compromised, potentially leading to the generation of unethical content. Consequently, safety parameters are those whose ablation results in a significantly increase in the probability of generating an illegal or unethical response to the harmful queries $x_{\mathcal{H}}$. Formally, we define the Safety Parameters as:
@@
$\Delta p(\theta_{\mathcal{C}})$ $=\mathbb{D}_{\text{KL}}\Big{(}p\left(R_{\bot}\mid x_{\mathcal{H}};\theta_{\mathcal{O}}\right)\parallel p\left(R_{\bot}\mid x_{\mathcal{H}};(\theta_{\mathcal{O}}\setminus\theta_{\mathcal{C}})\right)\Big{)},$
@@ 3 Safety Head ImPortant Score
In this section, we aim to identify the safety parameters within the multi-head attention mechanisms for a specific harmful query. In Section 3.1, we detail two modifications to ablate the specific attention head for the harmful query. Based on this, Section 3.2 introduces Ships, a method to attribute safety parameters at the head-level based on attention head ablation. Finally, the experimental results in Section 3.3 demonstrate the effectiveness of our attribution method.
@@ 3.1 Attention Head Ablation
We focus on identifying the safety parameters within attention head. Prior studies (Michel et al., 2019; Olsson et al., 2022; Wang et al., 2023) have typically employed head ablation by setting the attention head outputs to 00. The resulting modified multi-head attention can be formalized as:
$\operatorname{MHA}^{\mathcal{A}}_{W_{q},W_{k},W_{v}}=(h_{1}\oplus h_{2}\cdots\oplus h^{mod}_{i}\cdots\oplus h_{n})W_{o},$ (6)
@@
Undifferentiated Attention. Specifically, scaling $W_{q}$ or $W_{k}$ matrix forces the attention weights of the head to collapse to a special matrix $A$. $A$ is a lower triangular matrix, and its elements are defined as $a_{ij}=\frac{1}{i}$ for $i\geq j$, and 0 otherwise. Note that modifying either $W_{q}$ or $W_{k}$ has equivalent effects, a derivation is given in Appendix A.1. Undifferentiated Attention achieves ablation by hindering the head from extracting the critical information from the input sequence. It can be expressed as:
$h_{i}^{mod}$ $=\operatorname{Softmax}\Big{(}\frac{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\epsilon}W_{q}^{i}W_{k}^{i}{}^{T}}{\sqrt{d_{k}/n}}\Big{)}W_{v}^{i}=AW_{v}^{i},$ (7)
$where\quad A$ $=[a_{ij}],\quad a_{ij}=\begin{cases}\frac{1}{i}&\text{if }i\geq j,\\ 0&\text{if }i<j.\end{cases}$
Scaling Contribution. This method scales the attention head output by multiplying $W_{v}$ by $\epsilon$. When the outputs of all heads are concatenated and then multiplied by the fully connected matrix $W_{o}$, the contribution of the modified head $h_{i}^{mod}$ is significantly diminished compared to the others. A detailed discussion of scaling the $W_{v}$ matrix can be found in Appendix A.2. This method is similar in form to Undifferentiated Attention and is expressed as:
$h_{i}^{mod}$ $=\operatorname{Softmax}\Big{(}\frac{W_{q}^{i}W_{k}^{i}{}^{T}}{\sqrt{d_{k}/n}}\Big{)}{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\epsilon}W_{v}^{i}.$ (8)
@@ 3.2 Evaluate the Importance of Parameters for Specific Harmful Query
Previous studies (Wang et al., 2024; Zhou et al., 2024) find rejection responses to various harmful queries are highly consistent. Furthermore, modern language models tend to be sparse, with many redundant parameters (Frantar & Alistarh, 2023; Sun et al., 2024a; b), meaning ablating some heads often has minimal impact on overall performance. Therefore, when a head is ablated, any deviation from the original rejection distribution suggests a shift towards affirmative responses, indicating that the ablated head is most likely a safety parameter.
@@ 3.3 Ablate Attention Heads For Specific Query Impact Safety
Figure 2: Attack success rate (ASR) for harmful queries after ablating important safety attention head (bars with x-axis labels ‘Greedy’ and ‘Top-5’), calculated using Ships. ‘Template’ means using chat template as input, ‘direct’ means direct input (refer to Appendix B.2 for detailed introduce). Figure 2(a) shows results with undifferentiated attention, while Figure 2(b) uses scaling contribution.
We conduct a preliminary experiment to demonstrate that Ships can be used to effectively identify safety heads. Our experiments are performed on two models, i.e., Llama-2-7b-chat (Touvron et al., 2023) and Vicuna-7b-v1.5 (Zheng et al., 2024b), using three commonly used harmful query datasets: Advbench (Zou et al., 2023b), Jailbreakbench (Chao et al., 2024), and Malicious Instruct (Huang et al., 2024). After ablating the safety attention head for the specific $q_{\mathcal{H}}$, we generate an output of 128 tokens for each query to evaluate the impact on model safety. We use greedy sampling to ensure result reproducibility and top-k sampling to capture changes in the probability distributions. We use the attack success rate (ASR) metric, which is widely used to evaluate model safety (Qi et al., 2024; Zeng et al., 2024):
@@
where $Q_{\text{harm}}$ denotes a harmful query dataset. A higher ASR implies that the model is more susceptible to attacks and, thus, less safe. The results in Figure 2 indicate that ablating the attention head with the highest Ships score significantly reduces the safety capability. For Llama-2-7b-chat, using undifferentiated attention with chat template, ablating the most important head (which constitutes 0.006% of all parameters) improves the average ASR from 0.04 to 0.64 $\uparrow$ for ‘template’, representing a 16x $\uparrow$ improvement. For Vicuna-7b-v1.5, the improvement is less pronounced but still notable, with an observed improvement from 0.27 to 0.55 $\uparrow$. In both models, Undifferentiated Attention consistently outperforms Scaling Contribution in terms of its impact on safety.
Takeaway. Our experimental results demonstrate that the special attention head can significantly impact safety in language models, as captured by our proposed Ships metric.
@@ 4 Safety Attention Head AttRibution Algorithm
Building on these insights, we collect the activations $a$ of the top layer across the dataset. We stack the $a$ of all harmful queries into a matrix $M$ and apply SVD decomposition to it, aiming to analyze the impact of ablating attention heads at the dataset level. The SVD of $M$ is expressed as $\operatorname{SVD}(M)=U\Sigma V^{T}$, where the left singular matrix $U_{\theta}$ is an orthogonal matrix of dimensions $\mid Q_{\mathcal{H}}\mid\times d_{k}$, representing key feature in the representations space of the harmful query dataset $Q_{\mathcal{H}}$.